OpenStudy (anonymous):

What is the vertex for the graph of y = -2x^2 + 8x - 13? Answer (2, 11) (2, -5) (-2, -37) (-2, -21)

OpenStudy (anonymous):

I assume by 'vertex' you mean the turning point. By differentiation: \[\frac{d}{dx} -2x^2 + 8x - 13 = -4x + 8\] The derivative is 0 at turning points, so we get the equation \[ -4x + 8 = 0 \\ -4x = -8\\ x = 2 \]